# The iterated Dirichlet process and applications to Bayesian inference

### v1

@@ 6.1 IDPs on a finite state space
From (3.1) it follows that
${(\theta_{m0},\ldots,\theta_{m,L-1})\sim\operatorname{Dir}(\varepsilon p_{0},\ldots,\varepsilon p_{L-1}).}$
This gives
@@
We can apply this to $S^{\prime}=\{0,\ldots,L-1\}$, where $L\geq 2$. For each $\ell\in S^{\prime}=\{0,\ldots,L-1\}$, choose $B_{\ell}\in\mathcal{S}$ so that $\{B_{\ell}:\ell\in S^{\prime}\}$ is a partition of $S$. Define $\psi:S\to S^{\prime}$ by $\psi=\sum_{\ell=0}^{L-1}\ell 1_{B_{\ell}}$ and let $\xi^{\prime}=\{\xi^{\prime}_{ij}\}$ where $\xi^{\prime}_{ij}=\psi(\xi_{ij})$. Suppose we can only observe the process $\xi^{\prime}$. That is, we can only observe the values of $\xi$ with enough precision to tell which piece of the partition those values lie in. Based on some set of these observations, we wish to make probabilistic inferences about $\xi^{\prime}$. Since $\xi^{\prime}$ is an array of samples from an IDP on $S^{\prime}$, we may do this using the simplified formulas in this section.
@@ 7 H T T H H
One way to answer this is to model the example with an IDP as in Section 6.1. We take $L=2$, so that $S=\{0,1\}$, where 00 represents tails and $1$ represents heads. We then take $\kappa=\varepsilon=1$ and $p_{0}=p_{1}=1/2$. From the table above, we have $M=7$, $N_{m}=5$ for all $m$, and
@@ 6.3 Flicking thumbtacks
To model this example we take $L=2$, so that $S=\{0,1\}$, where 00 represents failure (point down) and $1$ represents success (point up). To match the modeling in [12], we take $\varepsilon=2$ and $p_{0}=p_{1}=1/2$, so that $\mathcal{D}(\varepsilon\varrho)\circ\pi_{1}^{-1}=\operatorname{Beta}(1,1)$, where $\pi_{1}:M_{1}\to[0,1]$ is the projection, $\nu\mapsto\nu(\{1\})$. We will use and compare two different values of $\kappa$ (which is denoted by $c$ in [12]). For the data, we have $M=320$ and $N_{m}=9$ for all $m$. Our row counts, $\overline{y}$, are given by $\overline{y}_{m1}=r_{m}$ and $\overline{y}_{m0}=9-r_{m}$.
@@ 6.4 Amazon reviews
The model in [12] only covers agents with two possible actions, such as coins and thumbtacks. The IDP, though, can handle agents whose range of possible actions is arbitrary.
@@ 6.5 The gamer distribution
[…]
for $x>0$. The fact that this is a probability density function is shown in Proposition 6.3 below. Note that if $X\sim\operatorname{Gamer}(r,c,\alpha)$ and $s>0$, then $sX\sim\operatorname{Gamer}(r,sc,\alpha)$.
@@
[…]
@@ 6.6.1 Players with matching scores
rank name hi score avg score IDP avg # games
@@ 10 The Matrix 15 13 43 2
The long-term average score of the player in the $m$th row of Table 3 will be $A(\theta_{m})$, where $A(x)=\sum_{\ell=0}^{499}\ell x_{\ell}$. For example, using (6.5), the expected long-term average score of Running Stardust is $E[A(\theta_{4})\mid Y=y]\approx 79.65$. These conditional expectations, rounded to the nearest integer, are shown in the “IDP avg” column of Table 4.
Looking at these averages, we can see at least two players whose numbers seem unusual. The first is Goat Radish. They played only one game and scored a 38, which is a relatively low score compared to the rest of the group. And yet the IDP model has given them an expected long-term average score of 71. Not only is this counterintuitive, it is also inconsistent with how the model treated The Pianist Spider.
@@
Our intuition is able to dismiss this line of reasoning because we know, for instance, that there is very little difference between a score of 38 and 39. Had Goat Radish scored a 39 instead, our predictions should not change that much. But we only know this because we are viewing the positive real numbers as more than just a set. We are viewing them as a totally ordered set with the Euclidean metric. The IDP model is not designed to utilize these properties of the state space. From its perspective, the number “38” is just a label. It is nothing more than the name of a particular element of the state space, and it happens to be an element that only two players were able to hit.
@@ 6.6.2 Matching scores removed
rank name hi score avg score IDP avg # games
@@ 6.6.3 Players with only a few games
rank name hi score avg score IDP avg # games
@@ 10 The Thing 40 24 26 10
Looking at Table 7, we see that Asparagus Soda corresponds to $m=9$ and Potato Log corresponds to $m=2$. Table 8 shows us that $E[A(\theta_{9})\mid Y=y]\approx 67$ and $E[A(\theta_{2})\mid Y=y]\approx 31$. In other words, the IDP model gives Asparagus Soda a much higher expected long-term average score than Potato Log. This confirms our intuition that Asparagus Soda is the better player. But because Asparagus Soda played only one game, the model should have a lot more uncertainty surrounding Asparagus Soda’s forecasted mean. To see this, we can compare approximate densities for $\mathcal{L}(A(\theta_{9})\mid Y=y)$ and $\mathcal{L}(A(\theta_{2})\mid Y=y)$. (See Figure 4.)

### v2

@@ 5.1 The case of a finite state space
From (2.1) it follows that $(\theta_{m0},\ldots,\theta_{m,L-1})\sim\operatorname{Dir}(\varepsilon p_{0},\ldots,\varepsilon p_{L-1})$. This gives
@@
We can apply this to $S^{\prime}=\{0,\ldots,L-1\}$, where $L\geq 2$. For each $\ell\in S^{\prime}=\{0,\ldots,L-1\}$, choose $B_{\ell}\in\mathcal{S}$ so that $\{B_{\ell}:\ell\in S^{\prime}\}$ is a partition of $S$. Define $\psi:S\to S^{\prime}$ by $\psi=\sum_{\ell=0}^{L-1}\ell 1_{B_{\ell}}$ and let $\xi^{\prime}=\{\xi^{\prime}_{ij}\}$ where $\xi^{\prime}_{ij}=\psi(\xi_{ij})$. Suppose we can only observe the process $\xi^{\prime}$. That is, we can only observe the values of $\xi$ with enough precision to tell which piece of the partition those values lie in. Based on some set of these observations, we wish to make probabilistic inferences about $\xi^{\prime}$. Since $\xi^{\prime}$ is an array of samples from an NDP on $S^{\prime}$, we may do this using the simplified formulas in this section.
@@ 7 H T T H H
One way to answer this is to model the example with an NDP as in Section 5.1. We take $L=2$, so that $S=\{0,1\}$, where 0 represents tails and $1$ represents heads. We then take $\kappa=\varepsilon=1$ and $p_{0}=p_{1}=1/2$. From the table above, we have $M=7$, $N_{m}=5$ for all $m$, and
@@ 5.3 Flicking thumbtacks
To model this example we take $L=2$, so that $S=\{0,1\}$, where 0 represents failure (point down) and $1$ represents success (point up). To match the modeling in [11], we take $\varepsilon=2$ and $p_{0}=p_{1}=1/2$, so that $\mathcal{D}(\varepsilon\varrho)\circ\pi_{1}^{-1}=\operatorname{Beta}(1,1)$, where $\pi_{1}:M_{1}\to[0,1]$ is the projection, $\nu\mapsto\nu(\{1\})$. We will use and compare two different values of $\kappa$ (which is denoted by $c$ in [11]). For the data, we have $M=320$ and $N_{m}=9$ for all $m$. Our row counts, $\overline{y}$, are given by $\overline{y}_{m1}=r_{m}$ and $\overline{y}_{m0}=9-r_{m}$.
@@ 5.4 Amazon reviews
The model in [11] only covers agents with two possible actions, such as coins and thumbtacks. The NDP, though, can handle agents whose range of possible actions is arbitrary.
@@ 5.5 The gamer distribution
$f(x)=\frac{rc^{r}}{\alpha^{r}\,\Gamma(\alpha)}\,x^{-r-1}\,\int_{0}^{\alpha x/c}y^{\alpha+r-1}e^{-y}\,dy,$ (5.7)
for $x>0$. The fact that this is a probability density function is a consequence of Proposition 5.3 below. Note that if $X\sim\operatorname{Gamer}(r,c,\alpha)$ and $s>0$, then $sX\sim\operatorname{Gamer}(r,sc,\alpha)$.
@@
According to Proposition 5.3, the parameter $r$ is the tail index of the mean player scores in the population. However, it is also the tail index of the raw player scores. To see this, let $\gamma(\beta,u)=\int_{0}^{u}y^{\beta-1}e^{-y}\,dy$ denote the lower incomplete gamma function. Then (5.7) can be rewritten as
@@ 5.6.1 Players with matching scores
rank name hi score avg score NDP avg # games
@@ 10 The Matrix 15 13 43 2
The long-term average score of the player in the $m$th row of Table 3 will be $A(\theta_{m})$, where $A(x)=\sum_{\ell=0}^{499}\ell x_{\ell}$. For example, using (5.5), the expected long-term average score of Running Stardust is $E[A(\theta_{4})\mid Y=y]\approx 79.65$. These conditional expectations, rounded to the nearest integer, are shown in the “NDP avg” column of Table 4.
Looking at these averages, we can see at least two players whose numbers seem unusual. The first is Goat Radish. They played only one game and scored a 38, which is a relatively low score compared to the rest of the group. And yet the NDP model has given them an expected long-term average score of 71. Not only is this counterintuitive, it is also inconsistent with how the model treated The Pianist Spider.
@@
Our intuition is able to dismiss this line of reasoning because we know, for instance, that there is very little difference between a score of 38 and 39. Had Goat Radish scored a 39 instead, our predictions should not change that much. But we only know this because we are viewing the positive real numbers as more than just a set. We are viewing them as a totally ordered set with the Euclidean metric. The NDP model is not designed to utilize these properties of the state space. From its perspective, the number “38” is just a label. It is nothing more than the name of a particular element of the state space, and it happens to be an element that only two players were able to hit.
@@ 5.6.2 Matching scores removed
rank name hi score avg score NDP avg # games
@@ 5.6.3 Players with only a few games
rank name hi score avg score NDP avg # games
@@ 10 The Thing 40 24 26 10
Looking at Table 7, we see that Asparagus Soda corresponds to $m=9$ and Potato Log corresponds to $m=2$. Table 8 shows us that $E[A(\theta_{9})\mid Y=y]\approx 67$ and $E[A(\theta_{2})\mid Y=y]\approx 31$. In other words, the NDP model gives Asparagus Soda a much higher expected long-term average score than Potato Log. This confirms our intuition that Asparagus Soda is the better player. But because Asparagus Soda played only one game, the model should have a lot more uncertainty surrounding Asparagus Soda’s forecasted mean. To see this, we can compare approximate densities for $\mathcal{L}(A(\theta_{9})\mid Y=y)$ and $\mathcal{L}(A(\theta_{2})\mid Y=y)$. (See Figure 4.)